10^{∘}N–25^{∘}N 0^{∘}–10^{∘}N Carib Total

MODIS

Annual bulk  143  35  37 215
Annual wet  23  10  9  42
DJF bulk  10  1  3  14
MAM bulk  27  21  14  62
JJA bulk  118  6  18 142
SON bulk  8  7  2  17

CMIP5

Annual bulk  44  11  10  65
Annual wet  17  9  4  30
DJF bulk  12  3  1  16
MAM bulk  10  4  2  16
JJA bulk  13  2  5  20
SON bulk  9  2  2  13